Circular No. 9124

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET C/2010 E1 (GARRADD) G. J. Garradd reports his discovery of a comet on CCD images taken

with the 0.5-m Uppsala Schmidt telescope at Siding Spring (discovery observation tabulated below), the object described as being more diffuse than nearby stars with a coma diameter of ∼ 15′′ and a possible faint tail ∼ 30′′ long in p.a. 250^{o} on a 40-s exposure. After posting on the Minor Planet Center’s ‘NEOCP’ webpage, numerous other CCD astrometrists have commented on the object’s cometary appearance. S. Foglia writes that thirty stacked 60-s exposures taken by R. Holmes (Ashmore, IL, U.S.A.) with a 0.81-m reflector on Mar. 12.13–12.14 UT show a coma of diameter 8′′–10′′ with a 60′′ tail in p.a. 230^{o}. Foglia adds that eight stacked 180-s exposures taken remotely by P. Concari, G. Galli, M. Tombelli, and himself with a 0.18-m reflector at the Tzec Maun Observatory (Moorook, Australia) on Mar. 12.44–12.46 also show a 10′′ coma. H. Sato (Ota-ku, Tokyo, Japan; remotely using a 0.25-m reflector at the RAS Observatory, Mayhill, NM, U.S.A.; Mar. 12.14–12.15) measured a 30′′ coma with a faint tail 45′′ long toward p.a. 230^{o}. E. Guido (Castellammare di Stabia, Italy) writes that he and G. Sostero also obtained images remotely with a 0.25-m reflector at the RAS Observatory near Mayhill, noting that fifteen co-added 120-s unfiltered exposures from Mar. 12.2 show a diffuse coma nearly 10′′ in diameter with a broad fan tail ∼ 20′′ long in p.a. 240^{o}. L. Elenin reports that six stacked 60-s unfiltered CCD exposures taken remotely with a 0.36-m astrograph at the Tzec Maun Observatory on Mar. 12.32 show a diffuse coma of size 0′.25-0′.3. P. M. Kilmartin measured a 1′ tail in p.a. 230^{o} on exposures taken by A. C. Gilmore and herself with the Mt. John 1.0-m reflector in poor seeing on Mar. 12.4. R. Ligustri (Talmassons, Udine, Italy) writes that twenty 60-s exposures taken remotely with a 0.40-m reflector at the RAS Observatory near Moorook on Mar. 12.7 show a coma of size 15′′ with a faint tail 15′′ long in p.a. 235^{o}.

2010 UT α_{2000} δ_{2000} Mag. Observer
Mar. 11.57146 9^{h}31^{m}37.^{s}23 −23^{o}35′45′′.1 17.4 Garradd

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2010-E59.

T = 2009 Oct. 24.604 TT  ω = 291.^{o}168
Ω = 169.281 }2000.0
q = 2.56144 AU  i = 71.721
2010 March 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green